# Minimal-doubling and single-Weyl Hamiltonians

### v1

@@ IV Conclusion and discussion
We then reconsidered BdG-based “single-Weyl” Hamiltonians within our minimally doubled framework. By introducing a Nambu particle/hole doubling and an appropriate species-splitting “Majorana” mass term into the Karsten-Wilczek minimal-doubling Hamiltonian, we showed that one Weyl node can be gapped out while the other remains gapless in the free theory. The protection mechanism is encoded in a momentum-dependent conserved generator acting in Nambu space: it commutes with the Hamiltonian, yet its spectrum is not quantized, which naturally suggests the Hamiltonian Ginsparg–Wilson–type structure. From this viewpoint, we studied a symmetry-preserving one-parameter deformation that keeps the full symmetry intact, yet can destabilize the single-node regime by generating additional Weyl nodes once the deformation exceeds a critical threshold. Concretely, the deformation opens a channel for extra gapless solutions that were gapped in the undeformed model, and the single-Weyl regime survives only inside a finite parameter window.
An important implication for the interacting theories is that the destabilizing deformation is not an artificial modification: it is not forbidden by the protecting symmetry, and accordingly, once gauge interactions are introduced, radiative corrections are expected to generate this term as an allowed counterterm with a renormalized coefficient that is not parametrically suppressed in lattice units, in close analogy with additive mass renormalization in Wilson fermion. The stability of the single-Weyl regime therefore could require a moderate tuning of the corresponding bare parameter so that the renormalized coefficient remains inside the single-node domain. In this sense, the exact non-onsite conservation alone does not guarantee radiative stability of the single-Weyl phase.

### v2

@@ IV Conclusion and discussion
We then reconsidered BdG-based “single-Weyl” Hamiltonians within our minimally doubled framework. By introducing a Nambu particle/hole doubling and an appropriate species-splitting “Majorana” mass term into the Karsten-Wilczek minimal-doubling Hamiltonian, we showed that one Weyl node can be gapped out while the other remains gapless in the free theory. The protection mechanism is encoded in a momentum-dependent conserved generator acting in Nambu space: it commutes with the Hamiltonian, yet its spectrum is not quantized, which naturally suggests the Hamiltonian Ginsparg-Wilson-type structure. From this viewpoint, we studied a symmetry-preserving one-parameter deformation that keeps the full symmetry intact, yet can destabilize the single-node regime by generating additional Weyl nodes once the deformation exceeds a critical threshold. Concretely, the deformation opens a channel for extra gapless solutions that were gapped in the undeformed model, and the single-Weyl regime survives only inside a finite parameter window.
An important implication for the interacting theories is that the destabilizing deformation is not an artificial modification: it is not forbidden by the exact non-onsite conservation law, and accordingly, once gauge interactions are introduced, radiative corrections are expected to generate this term as an allowed counterterm with a renormalized coefficient that is not parametrically suppressed in lattice units, in close analogy with additive mass renormalization in Wilson fermion. The stability of the single-Weyl regime therefore could require a moderate tuning of the corresponding bare parameter so that the renormalized coefficient remains inside the single-node domain. In this sense, the exact non-onsite conservation alone does not guarantee radiative stability of the single-Weyl phase.
Finally, let us comment on the relation of the present constructions to the Nielsen-Ninomiya no-go theorem. Under its standard assumptions, including locality, Hermiticity, translation invariance, and conserved onsite charges with discrete spectra, an isolated Weyl mode is forbidden. In the minimal-doubling constructions, the four-component Hamiltonian, which possesses two conserved onsite charges, yields two massless Dirac fermions, while the two-component Hamiltonian, with one conserved onsite charge, yields one massless Dirac fermion composed of two Weyl fermions. These constructions therefore remain within the standard scope of the theorem. By contrast, the BdG-based single-Weyl Hamiltonian possesses only a non-onsite momentum-dependent conserved generator, and hence lies outside that scope.
@@
From the condensed-matter viewpoint, single-Weyl Hamiltonians may serve as useful lattice models of Weyl-semimetal- or gapless-superconductor-type systems for studying how an isolated Weyl node can appear, how it can be destabilized by symmetry-allowed perturbations, and how additional nodes can emerge as parameters are varied.